OpenStudy (anonymous):

Why is the derivative of theta 1 and not 0. Isn't theta a number?

OpenStudy (unklerhaukus):

theta might be a variable

OpenStudy (anonymous):

Why is theta a variable? Wouldn't tan theta --> theta be a number in radians?

ganeshie8 (ganeshie8):

if \(\theta \) is a number/constant, \(\large \frac{d}{dx}(\theta) = 0\)

ganeshie8 (ganeshie8):

if \(\theta\) is a variable, \(\large \frac{d}{d\theta}(\theta) = 1\)

OpenStudy (anonymous):

So is theta a variable or number/constant?

Parth (parthkohli):

Theta is just like x. Is x a number? You decide.

OpenStudy (unklerhaukus):

theta usually represents an angle, if you are taking the derivative of the tangent of theta , you are trying to find the slope of that line for all possible angles theta

OpenStudy (agent0smith):

x is a number. y is a number. But they're variables, not constants.

OpenStudy (anonymous):

What is NDER to support computations of derivatives?